OpenStudy (anonymous):

1/2 x + 1/4 x = 16 16 21.3 32 64 i believe it is 21.3 someone check my answer?

OpenStudy (anonymous):

Spot on. Better to use 21 1/3 as its more accurate but yes. (1/2) x + (1/4) x = 16 2x + x = 64 3x = 64 x = 64/3 = 21 1/3